# When Your Boss Is an AI Bot: Exploring Opportunities and Risks of Manager Clone Agents in the Future Workplace

### 24 Sep 2025 v2

@@ 9. Discussion
This study examined how managers and workers envisioned the possibilities and risks of Manager Clone Agents in future workplaces. Drawing on workshop stories and discussions, we found that participants situated Manager Clone Agents at the intersection of productivity and relations: managers saw them as extensions of their authority and capacity, while workers emphasized how these systems would reshape recognition, fairness, and everyday labor. Across individual, interpersonal, and organizational levels, participants oscillated between hopes of enhanced productivity and fears of displacement or relationships erosion. Building on these findings, we contribute to ongoing debates in HCI and organizational research by situating Manager Clone Agents in three broader conversations: the distinctiveness of clone agents from other algorithmic management systems (see subsection 9.1), the rethinking of managerial work in the face of automation (see subsection 9.2), and the reconfiguration of worker–manager relations in an automated workplace (see subsection 9.3).
9.1. From Algorithmic Bosses to Cloned Managers: Reframing Managerial Authority in AI Systems
Much of the HCI and organizational literature on algorithmic management has focused on so-called algorithmic bosses, systems that allocate tasks, monitor workers, and enforce compliance in platform labor and other digitally mediated settings (Möhlmann et al., 2021; Jarrahi et al., 2020; Lee, 2016). These systems are typically characterized by impersonality: they operate through metrics, dashboards, and automated rules, with limited reference to the identities of specific human managers. Trust and legitimacy in such systems are often evaluated in terms of procedural fairness, transparency of rules, or predictability of outcomes (Lee et al., 2019b; Prassl, 2019).
Our findings extend this line of work by showing how Manager Clone Agents powered by GenAI, as a new form of AI management, are being re-imagined in fundamentally different terms. Unlike faceless algorithms, Manager Clone Agents are not abstract authorities but personalized proxies that explicitly embody a named individual. They carry the symbolic presence, voice, and persona of a particular manager, and in doing so transform how participants judged their legitimacy. For workers, what mattered was not only whether the system allocated tasks effectively and fairly, but whether it felt like an authentic extension of their manager. This personalization distinguishes Manager Clone Agents from prior models of algorithmic authority.
At the same time, the same symbolic link that granted Manager Clone Agents legitimacy could also undermine it. Both workers and managers expressed concern about the risks of misrepresentation, where an agent might speak or act in ways the human manager did not intend, thereby creating confusion over accountability. Participants stressed that while they valued clarification or reinforcement of existing managerial decisions, they were uneasy when Manager Clone Agents generated new responses or independent contributions, as this crossed into authorship that challenged identity and authority. These concerns highlight a paradox: the closer Manager Clone Agents come to cloning the manager, the higher the stakes when authenticity is uncertain.
This distinction suggests that Manager Clone Agents occupy a different conceptual space from algorithmic bosses or generic virtual assistants. Rather than replacing managerial oversight with impersonal rules, they aim to reproduce the relational authority of a specific human. This raises new questions for HCI and organizational theory about how authority is mediated, shared, or contested when a system is designed not simply to manage work, but to embody the manager. Our contribution here is to show that the design of AI systems in organizational contexts cannot be understood solely through the lens of efficiency, fairness, or control. With Manager Clone Agents, legitimacy and trust hinge on whether the technology can sustain the symbolic and relational bonds that management entails.
9.2. Rethinking Managerial Work in the Context of Automation
Research on managerial automation has often assumed a clear division between routine tasks that can be automated and higher-level strategic or relational work that should remain human-led (Shao et al., 2025; Zaidi et al., 2025; West, 2018). This division frames automation as a way of freeing up managers to focus on strategy, creativity, or leadership by removing administrative burdens (Raisch and Krakowski, 2021). Our findings complicate this assumption by showing that even routine managerial tasks, such as on-boarding new employees or providing quick feedback, carry symbolic and relational weight for workers. When delegated to Manager Clone Agents, they risk being reinterpreted as a lack of personal engagement, thereby weakening the interpersonal bonds that underpin managerial legitimacy.
Participants such as W3, W4, W8, M5 and M10 also imagined Manager Clone Agents as amplifiers of managerial capacity, as we showed in subsection 6.4, but again this was framed through relationships. By handling repetitive approvals or relaying pre-defined information, Manager Clone Agents were seen as allowing managers to invest more in external relationships and long-term vision. Yet participants insisted that the legitimacy of this amplification depended on boundaries. They stressed that workers looked to managers not only for efficiency, but also for discretion, authorship, and relational assurance that decisions came from a trusted human source. Delegating too far into the relational domain was perceived as hollowing out the leadership identity and undermining interpersonal trust between managers and teams.
These insights call for a rethinking of managerial work in the rise of GenAI. Management cannot be reduced to a bundle of discrete tasks, neatly divided into automatable and non-automatable categories. Instead, our findings emphasize its fundamentally relational character (Sluss and Ashforth, 2007). Managerial labor involves decision-making, but also symbolic and affective practices that maintain respect, recognition, and psychological safety within teams. For HCI, this shifts the challenge from designing automation that maximizes efficiency to designing systems that preserve the relational dimensions of leadership. By surfacing these nuances, our study contributes to a richer understanding of what aspects of management can be augmented by Manager Clone Agents and what must remain human to sustain the interpersonal bonds that give managerial work its meaning.
9.3. Rethinking Worker–Manager Relations in the Age of AI Agents
Another contribution of our study is to revisit how automation reshapes the relational fabric between workers and managers. Research on algorithmic management has often conceptualized this relationship as one of asymmetrical control, where digital systems strengthen managerial oversight while limiting worker agency (Barati and Ansari, 2022; Zhang et al., 2025; Xiao et al., 2025a). Our findings complicate this view by showing that Manager Clone Agents are imagined by our participants not only as tools of authority, but also as relational intermediaries that can redistribute opportunities, recognition, and burdens in uneven ways.
For the managers in our participants pool, Manager Clone Agents were valued as extensions of authority and capacity. They promised responsiveness in the manager’s absence, relief from repetitive administrative demands, and the ability to scale guidance to larger teams. Workers, however, approached Manager Clone Agents from a different vantage point: what mattered to them was whether the system preserved access to managers, created fair opportunities for visibility, and supported their career growth. While managers worried about identity and authorship, workers worried about recognition, legitimacy, and respect. These diverging stakes underline that automation in management cannot be understood solely through the lens of leadership efficiency but must account for its downstream effects on workers’ sense of belonging and advancement.
Our study contributes to a richer theorization of worker–manager relations in the age of AI agents. Rather than assuming automation uniformly empowers managers or uniformly disempowers workers, Manager Clone Agents highlight how authority, recognition, and trust are continually renegotiated in mediated interactions. For HCI and organizational research, this suggests that the legitimacy of AI in management hinges not only on technical reliability but on its ability to sustain relational equity between workers and managers.
9.4. Design and Policy Implications
Beyond theoretical contributions, our findings also highlight several implications for the design and governance of Manager Clone Agents. From a design perspective, Manager Clone Agents should incorporate mechanisms that clarify boundaries and preserve human agency. Participants valued augmentation but resisted substitution, suggesting that systems should make it explicit when an interaction is mediated by an agent, what kinds of contributions the agent is authorized to make, and how accountability is distributed. Design strategies could include graduated levels of initiative, ranging from lightweight reminders to bounded forms of delegation, so that managers and workers can calibrate the degree of automation appropriate to specific contexts. Equally important are features that enable workers to opt out, request human-only interactions, or signal when relational presence is essential, such as in mentoring or evaluation.
@@ 9.4. Design and Policy Implications
From a policy perspective, organizations adopting Manager Clone Agents must create safeguards that protect workers from new vulnerabilities. This includes ensuring that workers’ reliance on or resistance to Manager Clone Agents does not become a hidden basis for performance evaluation, and that managerial delegation does not erode opportunities for worker recognition and advancement. Policies should also mandate transparency about when and how Manager Clone Agents are deployed, as well as clear escalation paths for contesting errors or misrepresentations. At a broader level, regulatory frameworks may need to address the dual identity of Manager Clone Agents as both tools and proxies, clarifying where legal and ethical responsibility lies when agents act in a manager’s name.
@@ 10. Limitations and Future Work
Like any qualitative and speculative study, our work has several limitations that should be acknowledged. First, our findings are based on six design fiction workshops with 23 participants. While this approach enabled us to surface rich imaginaries and relational concerns, the sample is not representative of all industries, organizational sizes, or cultural contexts. Second, our use of speculative design elicited visions of possibilities and risks rather than documenting actual deployments. This allowed participants to reflect openly without being constrained by current technological limits, but it also means that their narratives may differ from how Manager Clone Agents are eventually integrated in practice. Third, our focus was primarily on managers and workers situated within formal organizations. Other stakeholders, such as HR professionals, policymakers, or AI developers, are likely to shape how Manager Clone Agents are designed and governed. Expanding the scope to include these perspectives would enrich understanding of the broader ecosystem in which Manager Clone Agents operate. at they enhance rather than erode the relational and cultural foundations of work.

### 29 Jan 2026 v3

@@ 8. Discussion
This study examined how managers and workers envisioned the possibilities and risks of Manager Clone Agents in future workplaces. Drawing on workshop stories and discussions, we found that participants situated Manager Clone Agents at the intersection of productivity and relations: managers saw them as extensions of their authority and capacity, while workers emphasized how these systems would reshape recognition, fairness, and everyday labor. Across individual, interpersonal, and organizational levels, participants oscillated between hopes of enhanced productivity and fears of displacement or relationships erosion. Building on these findings, we contribute to ongoing debates in HCI and organizational research by situating Manager Clone Agents in three broader conversations: the distinctiveness of clone agents from other algorithmic management systems (see subsection 8.1), the rethinking of managerial work in the face of automation (see subsection 8.2), and the reconfiguration of worker–manager relations in an automated workplace (see subsection 8.3). We then discussed the role of design fiction in exploring the future of work and future AI design (see subsection 8.4). We also offered design and policy implications towards fair, transparent and accountable AI intermediation (see subsection 8.5).
8.1. Why Manager Clone Agents Are Not Algorithmic Bosses: A New Mode of AI Authority
We define Manager Clone Agents as agentic AI systems designed to embody a specific human manager, including their communication style, tone, heuristics, and even symbolic presence (Liang, 2025; Lee et al., 2023; Cheng et al., 2025; The Journey Platform, 2025). Rather than acting as generic assistants, clone agents speak and perform managerial functions “as the manager,” stepping into meetings, clarifying decisions, or guiding workflows in the manager’s absence. Much of the HCI and organizational literature on algorithmic management focuses on algorithmic bosses—systems that assign tasks, monitor workers, and enforce compliance across platform labor and other digitally mediated work settings (Möhlmann et al., 2021; Jarrahi et al., 2020; Lee, 2016). These systems govern impersonally, relying on metrics, dashboards, and automated procedures. Within this framing, legitimacy is procedural: workers evaluate fairness, transparency, and predictability rather than who the system represents (Lee et al., 2019b; Prassl, 2019). By abstracting management away from individual managers, this literature implicitly treats authority as something that can be cleanly encoded in rules and extracted from social relationships. However, such abstractions overlook long-term forms of human authority that derive not from rules but from the social identities and interpersonal histories of specific managers; a gap that Manager Clone Agents make newly visible and bring into AI authority.
Firstly, the Manager Clone Agents re-intermediating algorithmic authority from the procedural domain into the relational and symbolic domain, where legitimacy hinges on social meaning rather than rule adherence. Participants did not judge Manager Clone Agents as faceless systems but as personalized proxies that stand in for a named leader. What mattered was whether it felt like an authentic extension of a specific manager. In our findings, AI-enabled management can be evaluated through relational criteria, for instance, whether the system carries the “voice,” moral sensibility, or interpersonal orientation of the manager. Clone agents expose an important limitation in existing theories of algorithmic management, which largely assume that algorithmic authority operates outside of social identity. By embodying the presence of a specific human, clone agents collapsed the conceptual boundary between algorithmic and human authority.
Secondly, this personalization introduced new vulnerabilities. The same symbolic closeness that granted clone agents legitimacy also raised the stakes when the agent deviated from expected behavior. Participants worried about moments when the agent generated responses the real manager would never give, or adopted tones and stances that felt incongruent. These deviations threatened the coherence of the manager’s identity, leading workers to question whether the AI was still “speaking as” the manager. As shown in previous literature about relational authority (Tyler and Lind, 1992; Wellman, 2017; Lake, 2009), when authority is relational, even small misalignments can destabilize trust in ways procedural errors do not.
Meanwhile, Clone agents also expanded managerial presence in new ways with AI agentic capability. Participants imagined the system representing managerial intent across time zones, onboarding new hires, mediating conflict, or explaining decisions based on inferred managerial logic. In effect, the agent made managerial identity computationally portable. Unlike algorithmic boss systems that impose abstract, depersonalized rules, these agentic capabilities computationally amplified the manager’s own authority and presence, extending what the human manager could accomplish.
Therefore our findings suggest that Manager Clone Agents occupy a conceptual space that existing models of algorithmic bosses do not capture. Legitimacy and trust are therefore tied to the agent’s ability to sustain the social meaning of managerial authority—not merely its ability to optimize workflows or enforce fairness. This reframing challenges HCI and CSCW scholars to expand theories of algorithmic management to account for systems that operate not as neutral infrastructures but as carriers of symbolic power, interpersonal history, and organizational meaning.
8.2. How Managerial Work and Its Legitimacy Are Reconfigured by Manager Clone Agents
Building on the previous section’s argument that Manager Clone Agents reshape the basis of managerial authority by shifting it from procedural to relational and symbolic grounds, this section turns to what these dynamics mean for the nature and future of managerial work itself. Rather than asking only whether clone agents can legitimately stand in for managers, our findings push us to reconsider a deeper question: what aspects of managerial labor give management its meaning, and how might automation reshape the legitimacy and identity of managerial work? By examining how participants interpreted delegation, amplification, and relational boundaries, we outline three conceptual shifts that challenge prevailing assumptions about what parts of management can or should be automated.
Firstly, our findings complicate the widely assumed distinction between routine managerial tasks suitable for automation and higher-level strategic or relational work that should remain human-led (Shao et al., 2025; Zaidi et al., 2025; West, 2018). Prior accounts imagine automation as lifting administrative burdens so managers can focus on more substantive leadership (Raisch and Krakowski, 2021). Yet participants repeatedly emphasized that even tasks that appear routine—such as onboarding a newcomer, acknowledging an effort, or sending a quick clarification—carry symbolic and relational significance. When performed by Manager Clone Agents, these actions can be reinterpreted as a withdrawal of personal engagement, thereby weakening the interpersonal recognition that workers rely on to evaluate managerial legitimacy. What appears functionally automatable may nonetheless remain relationally non-substitutable, revealing that managerial work cannot be cleanly partitioned into “automatable” and “non-automatable” categories without undermining its social meaning.
Secondly, although participants such as W7, M9, and M18 viewed Manager Clone Agents as promising amplifiers of managerial capacity for managers—as discussed in subsection 5.4—they emphasized that this amplification only maintained legitimacy when it respected relational boundaries. Clone agents were accepted when they handled repetitive approvals or relayed standardized updates, enabling managers to redirect their attention toward long-term vision or external relationships. Yet delegating too far into relational or affective domains was described as hollowing out the identity of leadership. This highlights that managerial work is fundamentally relational (Sluss and Ashforth, 2007), and managers must actively choose which interactions to keep as human-only encounters, knowing that these choices directly shape their perceived legitimacy and the integrity of their leadership identity.
Thirdly, our findings point toward a more speculative but increasingly plausible future in which the experience of being managed becomes a hybrid interactional ecology jointly mediated by humans and AI. Across workshops, participants imagined scenarios where workers interacted first, or even predominantly, with the clone agent rather than the human manager. If such patterns become normalized, managerial authority may depend less on managers’ in-person relational performances and more on their stewardship of the digital systems that now partially represent them. In such a future, managers are judged not only by how they lead directly, but by how responsibly they configure the tone, boundaries, and comportment of their AI counterparts. Legitimacy thus becomes distributed across the manager, the system’s reliability, and the manager’s governance of that system. In this context, managerial legitimacy shifts toward an infrastructural form: rooted in how managers design, maintain, and ethically govern the socio-technical environments through which workers experience leadership.
8.3. How AI Managerial Agents Reshape Labor, Power, and Worker Agency
Building on our earlier analysis of managerial authority, this section turns to the worker–manager relationship and asks how Manager Clone Agents reshape labor, power, and worker agency in AI-mediated organizations. Like in other labor studies literature (Barati and Ansari, 2022; Curchod et al., 2020; Monod, 2025), the relationship between managers and workers in our study is structured by deep and persistent power asymmetries in authority, evaluative control, access to information, and opportunities for visibility. Our findings consistently revealed these asymmetries in practice. Managers enjoyed privileged channels of recognition and decision-making: they delegated tasks, set priorities, interpreted performance, and decided when to appear or withdraw from everyday interactions. Workers, by contrast, relied on scarce moments of face-time to demonstrate competence, depended on managers to relay their concerns upward, and navigated the risks of being misunderstood, overlooked, or evaluated unfairly. Within this already asymmetric terrain, Manager Clone Agents do not arrive into a neutral field—they enter a workplace where one side holds disproportionate control over access, evaluation, and agenda-setting. Existing power inequalities still shape how clone agents are deployed, and how their effects are experienced.
However, parallel strands of research highlight that workers are not passive recipients of technological control. Workers actively develop tactics of resistance, appropriation, and reinterpretation in response to algorithmic systems (Ettlinger, 2018; Kellogg et al., 2020; Heemsbergen et al., 2022). Our findings also reveal a far more dynamic and contested terrain. Rather than a simple story of top-down authority, Manager Clone Agents complicate the relational fabric of work by redistributing visibility, recognition, and informational power in uneven and sometimes unexpected ways. To illuminate these tensions, we outline three shifts in how workers positively imagine their relationship to managers in a clone-mediated workplace.
First, Manager Clone Agents reshape the baseline power relation between workers and managers by redistributing how authority, access, and recognition circulate in everyday work. While managers in our study largely embraced clone agents as extensions of their authority and capacity, workers evaluated these same systems through a very different set of concerns. They focused on whether clone agents preserved meaningful access to managers, ensured fair visibility for their contributions, and supported rather than undermined their career opportunities.
Second, beyond these structural shifts, workers imagined immediate, everyday forms of agency emerging through their interactions with clone agents. They envisioned using agents to clarify ambiguous expectations, document contradictory instructions, and buffer emotionally difficult interactions. In this sense, clone agents become relational intermediaries that workers can strategically mobilize to protect themselves, secure clarity, or ensure commitments are honored. Such micro-practices illustrate how worker agency persists, and even expands, within AI-mediated management through subtle, everyday acts of reinterpretation and tactical use.
Third, participants articulated more radical future-oriented forms of counter-control, imagining scenarios where clone agents become instruments for challenging managerial opacity or renegotiating organizational power relations. In these speculative futures, workers leverage the agent’s persistent memory, consistency, and legibility to audit managerial behavior, hold managers accountable, or strategically navigate decision-making systems. Some imagined using clone agents to expose biases or inconsistencies; others envisioned querying the agent for rationales behind past decisions, thereby transforming managerial authority into something searchable and contestable. More radical imaginaries involved workers gaming the system to secure resources or reshape how managerial performance is evaluated. These visions extend everyday resistance into a broader reconfiguration of organizational power, suggesting that clone agents may enable workers to practice new forms of counter-control rooted in upward transparency and data-driven accountability (Kristiansen et al., 2018; Sun et al., 2023).
8.4. The Role of Design Fiction: Speculating the Agent’s Place in Future Work
In this study, design fiction served as more than a vehicle for imagining future technologies; it operated as a method for making future organizational possibilities thinkable (Wong, 2021; Ringfort-Felner et al., 2023). By introducing Manager Clone Agents into speculative workplace scenarios, design fiction enabled participants to articulate relational and power-based dynamics that are difficult to surface through traditional interview or observational methods.
Firstly, the fictional frame created a form of temporal and psychological distance that allowed participants to bracket present-day organizational constraints (Sloan, 2025). This distance enabled them to articulate forms of authority, recognition, and legitimacy that are normally taken for granted in everyday work. By stepping outside current norms, participants could ask questions that are not yet empirically available, such as how a computationally embodied manager might affect identity, trust, or the felt presence of leadership.
Secondly, design fiction opened a space for divergent and even contradictory future imaginaries (Tanenbaum et al., 2016). Rather than extending current algorithmic management trends, the speculative format elicited futures that ranged from incremental modifications to radically reorganized hierarchies. This plurality is essential: by engaging with futures that challenge or invert existing power structures, participants were able to articulate what must be protected (e.g., relational boundaries), what must be supported (e.g., worker agency), and what must be constrained (e.g., over-delegation of relational labor). Thus, design fiction not only surfaces multiple futures but also clarifies how managerial agents should be designed to remain accountable, relationally appropriate, and equitable across those futures.
Thirdly, design fiction functioned as an analytic probe (Schulte et al., 2016). The stories participants generated were not treated as predictions but as structured reflections that expose the relational, ethical, and organizational stakes embedded in emerging managerial agents. Through these narrative probes, we were able to observe how people reason about delegation limits, how they interpret the boundaries of managerial presence, and how they imagine exercising or retaining agency long before such agents materialize in practice.
8.5. Design and Policy Implications: Towards Fair, Transparent and Accountable AI Intermediation
Beyond theoretical contributions, our findings also highlight several implications for the design and governance of Manager Clone Agents. From a design perspective, Manager Clone Agents should incorporate mechanisms that clarify boundaries and preserve human agency. Participants valued augmentation but resisted substitution, suggesting that systems should make it explicit when an interaction is mediated by an agent, what kinds of contributions the agent is authorized to make, and how accountability is distributed. This requires visible and persistent disclosure cues, such as labeled message headers, distinctive visual avatars, or interaction summaries, that help workers distinguish when they are engaging with a clone versus the human manager. Our findings further indicate that workers evaluate legitimacy based on the perceived “domain” of the interaction; thus, interfaces should allow managers to designate certain contexts (e.g., onboarding, conflict navigation, performance feedback) as human-only zones and automatically prevent agent participation in those settings.
Design strategies could also include configurable tiers of initiative, allowing managers to set limits on whether the agent may proactively answer questions, surface insights, or merely relay prepared statements. For example, a clone agent might be restricted to providing clarifications in low-stakes administrative queries but require explicit manager approval before making recommendations in ambiguous or relationally sensitive situations. Workers in our study also emphasized the need for recourse: interfaces should provide quick mechanisms for requesting escalation to the human manager, annotating moments of misunderstanding, or pausing agent involvement when relational trust is at risk. Additionally, because many workers used clone agents as tools for emotional buffering, design should support private annotation spaces where workers can rehearse drafts, explore emotionally charged questions, or reflect before escalating to a human.
@@ 8.5. Design and Policy Implications: Towards Fair, Transparent and Accountable AI Intermediation
From a policy perspective, organizations adopting Manager Clone Agents must create safeguards that protect workers from new vulnerabilities. This includes ensuring that workers’ reliance on or resistance to Manager Clone Agents does not become a hidden basis for performance evaluation, and that managerial delegation does not erode opportunities for worker recognition and advancement. Policies should mandate enforceable worker controls: workers must be able to configure binding preferences that prevent agent participation in designated contexts (e.g., evaluation, promotion) without requiring managerial approval for each instance. Organizational guidelines must also require managers to retain responsibility for decisions made in their name, prohibiting the use of clone agents as shields for deflecting blame or obscuring authorship.
Transparency policies should mandate clear documentation of when clone agents are active, what training data they draw from, and what logs are retained, especially because workers in our study used agent memory to track fairness or resolve inconsistencies. Organizations should provide workers with controlled access to interaction histories involving clone agents so they can contest errors, ensure accurate representation, and prevent managerial misuse of selective memory.
At a broader level, regulatory frameworks may need to address the dual identity of Manager Clone Agents as both tools and proxies, clarifying where legal and ethical responsibility lies when agents act in a manager’s name. This could include requirements for audit trails that preserve the boundary between human intent and agent action, guidelines for preventing over-delegation of relational labor, and protections against reproducing or amplifying managerial biases in cloned decision patterns. Ultimately, governance must ensure that Manager Clone Agents enhance rather than erode relational equity: workers must have concrete mechanisms for contestation: the ability to flag and appeal agent decisions through channels that do not depend on managerial discretion, access to agent decision rationales, and formal representation in governance decisions about agent deployment.
@@ 9. Limitations and Future Work
Like any qualitative and speculative study, our work has several limitations that should be acknowledged. First, our findings are based on six design fiction workshops with 23 participants. While this approach enabled us to surface rich imaginaries and relational concerns, the sample is not representative of all industries, organizational sizes, or cultural contexts. Second, our use of speculative design elicited visions of possibilities and risks rather than documenting actual deployments. This allowed participants to reflect openly without being constrained by current technological limits, but it also means that their narratives may differ from how Manager Clone Agents are eventually integrated in practice. Third, our focus was primarily on managers and workers situated within formal organizations. Other stakeholders, such as HR professionals, policymakers, or AI developers, are likely to shape how Manager Clone Agents are designed and governed. Expanding the scope to include these perspectives would enrich understanding of the broader ecosystem in which Manager Clone Agents operate. Fourth, our study’s scope limits our ability to fully operationalize these principles into concrete artifact-level mechanisms without deeper engagement with workers’ lived experiences of power asymmetries. Similarly, the difficulty participants faced in imagining structural alternatives to existing hierarchies within mixed manager-worker workshops is notable. Future work could explicitly target these gaps through workers-only speculative sessions and participatory design focused on the practical mechanics of contestation and control.